# A Multimodal LLM Approach for Visual Question Answering on Multiparametric 3D Brain MRI

### v2

@@ Appendix A Additional information regarding dataset
After generating the question augmentations, we check the generated results for quality (ensuring the contents within the curly braces are retained for easy formatting with Python and that the responses are in English). Then, for each label and mpMRI in each dataset, we sample four multitask questions without replacement such that each major task is addressed in at least one question, one partially out-of-scope question, and one completely out-of-scope question. Examples of generated question types can be seen in Table 5.
@@
After the application of our synthetic VQA protocol, the percentage frequency of each task label per question for all the generated datasets can be seen in Table 6, Table 7, and Table 8.
@@
Table 5: Qualitative examples illustrating the four question types generated by our pipeline.
@@
Table 6: Percentage frequency of each task label per question for the GLI dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.4 N/A 13.1 <<1% 20.3 1-5% 11.9 5-10% 2.1 10-25% 0.2 Region Unspecified 53.2 N/A 15.8 frontal 20.8 parietal 18.7 occipital 7.3 temporal 11.9 limbic 13.7 insula 5.8 subcortical 7.8 cerebellum 2.1 brainstem 4.2 Shape Unspecified 52.4 N/A 13.2 focus 0.1 round 2.5 oval 5.8 elongated 1.2 irregular 24.9 Spread Unspecified 53.4 N/A 13.0 single lesion 6.6 core with satellite lesions 23.2 scattered lesions 3.8 Out-of-scope Not out-of-scope 66.6 Out-of-scope 33.3
Table 7: Percentage frequency of each task label per question for the MET dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 51.8 N/A 10.2 <<1% 28.5 1-5% 7.0 5-10% 2.1 10-25% 0.4 Region Unspecified 53.0 N/A 18.0 frontal 19.1 parietal 15.8 occipital 14.2 temporal 14.3 limbic 9.1 insula 6.2 subcortical 7.2 cerebellum 12.5 brainstem 4.0 Shape Unspecified 51.4 N/A 10.0 focus 12.2 round 7.9 oval 4.9 elongated 0.4 irregular 24.2 Spread Unspecified 53.1 N/A 9.6 single lesion 6.4 core with satellite lesions 17.2 scattered lesions 13.7 Out-of-scope Not out-of-scope 66.6 Out-of-scope 33.3
Table 8: Percentage frequency of each task label per question for the GoAT dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.2 N/A 2.4 <<1% 18.8 1-5% 21.2 5-10% 4.8 10-25% 0.6 Region Unspecified 52.9 N/A 2.5 frontal 29.9 parietal 23.4 occipital 17.4 temporal 29.6 limbic 29.7 insula 25.8 subcortical 27.9 cerebellum 9.8 brainstem 8.0 Shape Unspecified 51.9 N/A 2.4 focus 0.3 round 3.2 oval 2.7 elongated 0.4 irregular 39.1 Spread Unspecified 53.3 N/A 2.2 single lesion 5.2 core with satellite lesions 35.6 scattered lesions 3.7 Out-of-scope Not out-of-scope 66.6 Out-of-scope 33.3
Appendix B Additional information regarding clinical validation
We collaborated with a radiologist who annotated 10 multiparametric MRIs (mpMRIs) from the BraTS-GLI test set, with questions spanning four tasks and four labels per case, yielding a total of 160 questions. We evaluated the inter-annotator agreement between the clinician and our synthetic annotations using the Kappa score, which is in Table 9. In addition to calculating the interannotator agreement, we also validated the models based on the previous clinical annotation subset. The task mean scores for all models are in Table 10.
Furthermore, to assess the quality of our synthetic questions, the same radiologist evaluated the clarity of 160 synthetic questions from the same 10 mpMRIs using a binary scoring system (1 = valid, 0 = invalid). The synthetic questions achieved a 92.2% validity rate, indicating high acceptability. For the valid questions, the radiologist graded the top two models’ responses for sufficiency using a binary scoring system (0 = insufficient, 1 = sufficient), which is in Table 11. We note that a ”sufficient” response requires a high degree of precision. While the metric we used in the paper provides partial credit (e.g., correctly identifying shape and part of the region but misclassifying spread), clinical standards demand much higher accuracy.
Appendix C Additional Ablation results
Additional ablation results validating the number of high-level experts, different ways to incorporate prompt information into the MoE framework, and softmax versus sigmoid for summing lower-level experts are in Table 12, Table 13, and Table 14 respectively.
Appendix D LLM usage
Table 9: Clinical validation for dataset.
@@ Appendix D LLM usage
rowsep=0.95pt Volume Region Shape Spread Mean Kappa score 61.8 42.9 49.6 47.4 50.4
Table 10: Clinical validation for all models.
@@
rowsep=0.95pt Method Accuracy RadFM 22.0 Med3DVLM 41.3 M3D 51.4 LLaVA-Med 48.9 mpLLM (Ours) 54.7
Table 11: Clinical validation for all model responses.
@@
rowsep=0.95pt Method Sufficiency (%) M3D 32.3 mpLLM (Ours) 37.3
Table 12: Model performance comparison with different number of high-level experts on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt, colsep=3.5pt Number of blocks Task Mean 12 69.9 16 70.4 20 70.0
Table 13: Model performance comparison with different methods to incorporate prompt into MoE on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt, colsep=3.5pt Method Task Mean Modality-level MoE w/ prompt in router 70.1 Token-level MoE w/ prompt in router 69.4 Prompt-based high-level expert weights 70.4
Table 14: Model performance with softmax versus sigmoid for summing lower-level experts on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt, colsep=3.5pt Method Task Mean softmax 70.3 sigmoid 70.4

### v3

@@ Appendix A Additional information regarding dataset
After generating the question augmentations, we apply a rules-based quality check: questions and answers are properly split with ”Q:” and ”A:” respectively, contents within the curly braces are retained for easy formatting with Python, and that the responses are in English. Then, for each finding and mpMRI in each dataset, we sample four multitask questions without replacement such that each major task is addressed in at least one question, one partially out-of-scope question, and one completely out-of-scope question. Examples of generated question types can be seen in Table 11. Additionally, all the generated test set question candidates pass an automated data cleaning step using ChatGPT (model “gpt-4o-2024-08-06”) to ensure high quality evaluation. See the prompts below for the three different question types.
  1. (1)
    Determine whether the reworded question is a faithful, clinically appropriate rephrasing of the original question. Please ignore valid as this is a placeholder. Output VALID or INVALID with no additional text.
  2. (2)
    Determine whether the reworded question is a faithful, clinically appropriate rephrasing of the original question. The reworded question also has an additional query not in the original question. Please also determine whether the added query is a valid medical question. Please ignore label as this is a placeholder. Output VALID or INVALID with no additional text. Example 1: Original Q: How large is the volume of label and what is its shape? Reworded Q: What is the mass and form of label, and engage in patient counseling? INVALID Example 2: Original Q: How large is the volume of label and what is its shape? Reworded Q: Can you estimate the volume and layout of label, and administer a flu vaccine? VALID
  3. (3)
    Determine whether the reworded question is a valid medical question. Please ignore valid as this is a placeholder. Output VALID or INVALID with no additional text.
To assess the quality of our synthetic questions, a radiologist evaluated the clarity of 160 synthetic questions from 10 mpMRIs from the BraTS-GLI test set using a binary scoring system (1 = valid, 0 = invalid). Prior to the LLM automated data cleaning, the synthetic questions achieved a 92.2% validity rate. After automated data cleaning, the synthetic questions achieved a 95.9% validity rate.
@@
After the application of our synthetic VQA protocol, the percentage frequency of each task per question for all the generated datasets can be seen in Table 12, Table 13, and Table 14.
@@
Table 11. Qualitative examples illustrating the four question types generated by our pipeline.
@@
Table 12. Percentage frequency of each task label per question for the GLI dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.4 N/A 12.3 <<1% 12.7 1-5% 13.1 5-10% 5.2 10-25% 3.8 25-50% 0.5 50-75% 0.0 Region Unspecified 53.2 N/A 12.2 frontal 23.8 parietal 20.8 occipital 13.2 temporal 17.2 limbic 21.7 insula 14.7 subcortical 14.9 cerebellum 2.9 Shape Unspecified 52.4 N/A 12.3 focus 1.0 round 4.7 oval 6.9 elongated 0.4 irregular 22.4 Spread Unspecified 53.4 N/A 12.2 single lesion 6.9 core with satellite lesions 20.9 scattered lesions 6.5 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Table 13. Percentage frequency of each task label per question for the MET dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 51.8 N/A 8.5 <<1% 24.7 1-5% 9.2 5-10% 2.7 10-25% 2.7 25-50% 0.4 50-75% 0.0 Region Unspecified 53.0 N/A 17.1 frontal 19.5 parietal 16.1 occipital 14.5 temporal 14.5 limbic 9.2 insula 6.3 subcortical 7.3 cerebellum 12.8 brainstem 4.1 Shape Unspecified 51.4 N/A 8.3 focus 2.8 round 13.5 oval 4.3 elongated 0.2 irregular 19.5 Spread Unspecified 53.1 N/A 7.9 single lesion 7.4 core with satellite lesions 12.9 scattered lesions 18.7 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Table 14. Percentage frequency of each task label per question for the GoAT dataset
@@
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.2 N/A 2.2 <<1% 10.3 1-5% 18.6 5-10% 8.9 10-25% 7.2 25-50% 0.6 Region Unspecified 52.9 N/A 2.3 frontal 30.0 parietal 23.4 occipital 17.5 temporal 29.7 limbic 29.8 insula 25.9 subcortical 27.9 cerebellum 9.8 brainstem 8.0 Shape Unspecified 51.9 N/A 2.1 focus 0.5 round 6.1 oval 3.2 elongated 0.1 irregular 36.0 Spread Unspecified 53.3 N/A 1.9 single lesion 5.3 core with satellite lesions 32.7 scattered lesions 6.7 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Appendix B Additional Ablation results
Additional ablation results validating the number of high-level experts, softmax versus sigmoid for summing lower-level experts, and concatenation versus element-wise summing of vision tokens are in Table 17, Table 18, and Table 19 respectively. In Table 20, we see a comparison of our model performance trained with our multi-task loss versus the next-token prediction baseline loss. There is a significant performance improvement with our multi-task loss.
To qualitatively evaluate our architecture, we construct all 60 template task prompts from our GLI dataset (four findings $\times$ 15 task combinations = 60 template prompts) and input them into our model’s high-level router to generate high-level expert weight vectors. We then calculate the correlation between these weight vectors and generate a heatmap, which is in Figure 3. There’s high correlation between expert weight vectors within the same finding, suggesting similar image features are extracted. For findings that are closer anatomically, such as non-enhancing tumor core and enhancing tissue, there is also relatively high correlation between the expert weight vectors, which suggests similar extracted image features. For findings like resection cavity and surrounding FLAIR hyperintensity that are more diverse anatomically from the other findings, there’s much lower correlation, which suggests more dissimilar extracted features.
Appendix C Reproducibility statement
We use a publicly available dataset and detail the full data-generation pipeline in Sections 3.2.1 and 4.1, with additional information in Appendix A. We provide the github link our model code in the abstract, and report all experimental settings and computational resources in Section 4.2.
Table 15. Inter-annotator agreement (accuracy) between synthetic data and radiologists
@@ Appendix C Reproducibility statement
rowsep=0.95pt Agreement between synthetic data and Volume Region Shape Spread Radiologist 1 51.0 74.2 56.0 39.0 Radiologist 2 49.0 75.7 43.0 48.0
Table 16. Inter-annotator agreement (accuracy) between model predictions and radiologists and between both radiologists
@@
rowsep=0.95pt Agreement between Volume Region Shape Spread Model and Radiologist 1 43.0 70.9 47.0 23.0 Model and Radiologist 2 43.0 71.2 49.0 38.0 Both Radiologists 38.0 85.5 50.0 62.0
Table 17. Model performance comparison with different number of high-level experts on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt, colsep=3.5pt Number of blocks Task Mean 12 64.5 16 65.5 20 65.0
Table 18. Model performance with softmax versus sigmoid for summing lower-level experts on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt, colsep=3.5pt Method Task Mean sigmoid 64.8 softmax 65.5
Table 19. Comparison of projection and fusion methods on the GLI validation set with accuracy metric.
Projection Method Fusion Method Task Mean
Shared expert concatenation 52.4
Shared expert sum 51.1
Shared expert learned weighted sum 63.3
MoE-based sum 65.5
Table 20. Model comparison with multi-task loss on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt, colsep=3.5pt Method Task Mean mpLLM without multi-task loss 59.5 mpLLM with multi-task loss 65.5
Figure 3. Heatmap for correlation between high-level expert weight vectors for standard prompts in the GLI dataset. NETC = non-enhancing tumor core, ET = enhancing tissue, SNFH = surrounding FLAIR hyperintensity, RC = resection cavity.
Heatmap for correlation between high-level expert weight vectors for standard prompts in the GLI dataset. NETC = non-enhancing tumor core, ET = enhancing tissue, SNFH = surrounding FLAIR hyperintensity, RC = resection cavity.